OpenStudy (anonymous):

In right triangle ABC shown to the right, if AB=14 and BC=12, then BD=

OpenStudy (anonymous):

sqrt(1002/7)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

Cos B = 6/7 so BD = 12 Cos B

OpenStudy (anonymous):

what comes after that?

OpenStudy (anonymous):

and how did you get 6/7?

OpenStudy (anonymous):

Can you not multiply 12 by 6/7? (and 6/7 = 12/14)

OpenStudy (anonymous):

i don't quite understand the method in which you are using the find the answer; could you elaborate a little more?

OpenStudy (anonymous):

Cos = Adjacent/Hypotenuse